# A Purely Data-Driven Adaptive Impedance Matching Method Robust to Parasitic Effects

### v1

@@ II-C Impedance Matching with Parasitic Effects
Considering the inherent limitations of the manufacturing process, parasitic parameters may be introduced at any location along the RF channel connecting the antenna to the transmitter’s RFFE (e.g., radio frequency integrated circuits or power amplifier modules). As shown in Fig. 2, parasitic effects in RF channel arise from three key aspects.
  • L-network itself parasitics: These parasitic effects primarily stem from the manufacturing process of the TMN itself [34]. Specifically, whether employing continuously tunable components (e.g., varactors [35] or MEMS varactors [36]) or digitally tunable components (e.g., MEMS switches [37], PIN diodes [12], or CMOS switches [38]), each may introduce additional parasitic capacitance, inductance and resistance. In the ideal case, the L-network transforms the load impedance $Z_{L}$ into the input impedance $Z_{\text{in}}$ via the function $\mathcal{H}(\cdot)$, as described by Eq. (1). The presence of parasitic effects modifies this transformation to a new mapping $\mathcal{H}^{\prime}(\cdot)$, which maps the intermediate impedance $Z_{1}$ to $Z_{2}=\mathcal{H}^{{}^{\prime}}(Z_{1};f)$, where $f$ denotes the operating frequency.
  • RFFE interface parasitics: These parasitic effects arise from the physical interconnection between the TMN and the RFFE. During the PCB routing, via formation and soldering processes [39], connection discontinuities or non-ideal contacts can introduce additional parasitic inductance, capacitance, and resistance. The impedance transformation resulting from these parasitic effects is denoted by $\mathcal{H}_{2}(\cdot)$, which maps the intermediate impedance $Z_{2}$ to the final input impedance $Z^{{}^{\prime}}_{in}=\mathcal{H}_{2}(Z_{2};f)$.
@@
Above all, due to the presence of parasitic effects, the impedance matching condition for ideal L-nerwork given in Eq. (4) is modified to
@@
Neglecting these parasitic effects will lead to impedance matching failure. Taking parasitic effects into account, the overall impedance transformation is expressed as $(\mathcal{H}_{2}\circ\mathcal{H}^{\prime}\circ\mathcal{H}_{1})(\cdot)$, which is strongly nonlinear and inherently uncertain. The strong nonlinearity arises from the fact that parasitic elements are typically superimposed on RF channel in a mixture of series and shunt configurations. This complex interconnection results in a high-order rational function form for the input impedance. Furthermore, the inherent uncertainty stems from the unpredictable nature of parasitic parameters. Specifically, the types, values and positions of parasitic parameters are unknown, making the exact mathematical form of $\mathcal{H}_{1}(\cdot)$, $\mathcal{H}_{2}(\cdot)$ and $\mathcal{H}^{\prime}(\cdot)$ indeterminate. Consequently, the presence of parasitic effects not only complicates the relationship between the tunable components and the input impedance but also introduces inherent uncertainty, thereby significantly increasing the difficulty of adaptive impedance matching.
@@ III Behavioral Modeling of Circuits with Parasitic Effects
In this section, we will first employ multiport network theory to demonstrate that S-parameters can effectively characterize the circuit behavior of TMN in the presence of parasitic effects. Subsequently, we will introduce RECBM-Net, a deep learning model designed to fit the TMN behavior.
@@ III-A Analyzing TMN Using MultiPort Network Theory
Parasitic effects cause the practical TMN to deviate from the ideal L-network, resulting in a complex and uncertain topology. Therefore, it is not feasible to characterize TMN behavior from the perspective of impedance transformation. By modeling the practical TMN as a two-port network and analyzing it using multiport network theory [40], we can explore the circuit behavior of TMN under parasitic effects without requiring detailed knowledge of internal structure.
Any passive two-port network representing a practical TMN can be equivalently represented as a cascade of series and shunt arms. As illustrated in Fig. 3, the equivalent cascaded network of a practical L-network comprises $M$ series arms and $N$ shunt arms due to the parasitic effects. The transmission matrix (ABCD-matrix) of a two-port network with a series arm is given by
@@
If the first arm is shunt (rather than series), the first matrix on the right side of Eq. (9) is set to the identity matrix. Similarly, if the last arm is series (rather than shunt), the last matrix is set to the identity matrix. Consequently, the ABCD matrix of any passive two-port network can be expressed in the general form of Eq. (9), regardless of the specific arrangement of series and shunt arms. Finally, based on the conversion relationship between ABCD matrix and scattering parameters (S-parameters), the S-parameters of the entire two-port network is given by
@@
where $Z_{0}$ represents the characteristic impedance. The S-parameters describe the reflection and transmission characteristics of $n$-ports networks in high-frequency. They effectively characterize the circuit behavior of practical TMN in the presence of parasitic effects. In addition, the S-parameters depend exclusively on the network’s internal topology and are not influenced by the load impedance.
Equations (9) and (10) establish the S-parameters of a practical TMN as a function of the operating frequency $f$ and the tunable capacitors $C_{p}$ and $C_{s}$. We donate this mapping as
@@ III-B DNN based modeling
$\mathcal{F}$ is an extremely complex and highly nonlinear function. According to the universal approximation theorem [41], a multi-layer feedforward network with a sufficient number of hidden neurons can approximate any continuous function with arbitrary precision. By training a DNN on the measured S-parameters of the practical TMN, we can obtain an accurate estimate $\hat{\mathcal{F}}$ of the mapping $\mathcal{F}$. The preceding analysis of the fundamental mathematical representation of $\mathcal{F}$ is crucial, as it confirms the existence of a mathematical relationship between $(f,C_{p},C_{s})$ and the S-parameters. Leveraging a purely data-driven approach, DNN directly establishes an end-to-end nonlinear mapping. This method fundamentally circumvents the challenge of determining detailed topology while enabling accurate characterization of system behavior under parasitic effects.
The structure of the proposed RECBM-Net is illustrated in Fig. 4. It consists of an input module, a TMN operating state encoder, a circuit behavior decoder, an output mudule and a residual structure. The input vector is defined as $\mathbf{x}=(f,C_{p},C_{s})\in\mathbb{R}^{3}$. The output of the RECBM-Net is an 8-dimensional vector $\mathbf{y}\in\mathbb{R}^{8}$, where each pair of entries corresponds the real and imaginary components of one of the S-parameters, i.e., $\Re(S_{11})$, $\Im(S_{11})$, $\Re(S_{12})$, $\Im(S_{12})$, $\Re(S_{21})$, $\Im(S_{21})$, $\Re(S_{22})$ and $\Im(S_{22})$. $\Re(\cdot)$ and $\Im(\cdot)$ denote the real and imaginary parts, respectively. The proposed RECBM-Net employs an encoder-decoder structure with nine hidden layers. In operating state encoder module, the hidden layer dimensionality is initially increased to capture deeper and more abstract features from the input. Specifically, the network progressively expands the feature space with 64, 128, 256, 512, and 1024 neurons. Through the operating state encoder, the 3-dimensional TMN operating state $(f,C_{p},C_{s})$ is transformed into a 1024-dimensional representation, thereby enabling RECBM-Net to capture the intricate nonlinear relationships between the operating state and the corresponding S-parameters. Subsequently, in the circuit behavior decoder module, the network begins to reduce the dimensionality in the later layers with 512, 256, 128, and 64 neurons, effectively decoding the learned features into a compact representation. The final output layer consists of 8 neurons, directly predicting the real and imaginary parts of the S-parameters.
@@ IV Matching Solution Determination Strategy Based on Surrogate Model
In this section, we first formulate the task of determining matching solution based on a circuit behavior surrogate model as an mathematical optimization problem. Given the difficulty of solving this problem analytically, we will propose three optimization strategies with varying computational overhead.
@@ IV-A Formulating the Matching Optimization Problem
Figure 5: Overview of the proposed purely data-driven adaptive impedance matching method. The RECBM-Net is first trained using measured S-parameters of the practical L-network. During operation, it predicts the $\hat{\mathbf{S}}_{\text{now}}$ corresponding to the current TMN operating state. According to Eq. (12), the impedance sensor then predicts the $\hat{\Gamma}_{L}$. Finally, the control unit determines the matching solution based on $f$ and $\hat{\Gamma}_{L}$. To achieve this, we propose three strategies that differ in the computational overhead.
@@
where $\psi(\cdot)$ denotes taking the magnitude of the reflection coefficient given in Eq. (13). Additionally, the forward propagation of RECBM-Net for predicting $\mathbf{\hat{S}}$ is denoted as ^$\hat{\mathcal{F}}(f,C_{p},C_{s})$. Based on this black-box model of TMN behavior, the impedance matching process can be fomulated as the following mathematical optimization problem:
@@
where $C_{p,\min}$ and $C_{p,\max}$ denote the lower and upper limits of $C_{p}$, while $C_{s,\min}$ and $C_{s,\max}$ donate those of $C_{s}$. In the current mismatch scenario, the operating frequency $f$ and the derived load reflection coefficient $\hat{\Gamma}_{L}$ are treated as fixed.
Input: Number of particles $N$, individual learning factor $\kappa_{1}$, social learning factor $\kappa_{2}$, cooling factor $\lambda$, maximum iterations $\mathcal{I}_{\max}$, load reflection $\hat{\Gamma}_{\text{L}}$, frequency $f$, threshold $\varepsilon$, tuning range $\mathcal{C}=[C_{p,\min},C_{p,\max}]\times[C_{s,\min},C_{s,\max}]$.
@@
19             $\mathbf{v}_{n}^{(t)}&=\chi[\mathbf{v}_{n}^{(t-1)}+\kappa_{1}r_{1}(\mathbf{P}_{\text{best},n}^{(t-1)}-\mathbf{x}_{n}^{(t-1)})\\ &\quad+\kappa_{2}r_{2}(\mathbf{G}_{\text{rand}}^{(t-1)}-\mathbf{x}_{n}^{(t-1)})].\end{aligned}$
@@
22                   $\Gamma_{n}^{(t)}=1000$.
@@
27             Updated: $\mathbf{P}_{\text{best},n}^{(t)}$, $\Gamma_{\text{best},n}^{(t)}$,$\mathbf{G}_{\text{best}}^{(t)}$ and $\Gamma_{\text{best}}$.
@@
2Initialization: $t=0$, $\boldsymbol{\theta}^{(0)}=\boldsymbol{\theta}_{0}$, $\mathbf{m}_{0}=0$,$\mathbf{v}_{0}=0$.
@@
10       Evaluate reflection coefficient: $\Gamma_{\text{in}}^{(t)}=\psi(\hat{\mathcal{F}}(f,\boldsymbol{\theta}^{(t)}),\hat{\Gamma}_{L})$.
@@ IV-B Strategies for Solving the Matching Problem
Given that the RECBM-Net comprises multiple hidden layers with numerous neurons and activation functions, $\hat{\mathcal{F}}(\cdot)$ is an extremely complex nonlinear function. Embedding $\hat{\mathcal{F}}(\cdot)$ within the objective function $\psi(\cdot)$ further complicates the derivation of an analytical optimal solution. To address this, we first introduce two classic numerical optimization strategies: SAPSO and AD-Adam. These methods require varying numbers evaluations of RECBM-Net to compute the reflection coefficient for each TMN configuration, leading to considerable online computational overhead. To eliminate the need for repeated RECBM-Net inference during optimization, we further propose an inverse mapping solver network (IMS-Net) for RECBM-Net to directly obtain the matching solution.
Classic strategy 1 (SAPSO): SAPSO combines the fast convergence and high accuracy of particle swarm optimization (PSO) with the ability of simulated annealing (SA) to escape local optima by accepting occasionally worse solutions. This hybridization enhances the global search for an optimal TMN configuration. Using SAPSO to obtain the matching solution involves three major steps.
  • Initialization of the population: Let $n$ indexes particles, and $t$ indexes iterations. At iteration $t=0$, set each particle’s initial position $\mathbf{x}_{n}^{(0)}$ (representing a candidate TMN configuration) and velocity $\mathbf{v}_{n}^{(0)}$ within the given constraints. For each particle $n$, compute the initial fitness $\Gamma_{n}^{(0)}$ and record both its personal best position $\mathbf{P}_{\text{best},n}^{(0)}$ and the global best position $\mathbf{G}_{\text{best}}^{(0)}$ among all particles. Based on $\mathbf{G}_{\text{best}}^{(0)}$, determine the initial temperature $T$ for the simulated annealing component. Additionally, initialize the compression factor $\chi$ and the annealing rate $\delta$.
  • Velocity and position updates: At iteration $t$, first calculate the normalized Metropolis acceptance probabilities $p_{n}^{(t-1)}$ for each particle $n$ based on its personal best fitness $\Gamma_{\text{best},n}^{(t-1)}$. $\Gamma_{\text{best},n}^{(t-1)}$ donates the best fitness value achieved by particle $n$ up to the current iteration. Next, perform a roulette wheel selection using $p_{n}^{(t-1)}$ to obtain a candidate global best position $\mathbf{G}_{\text{rand}}^{(t-1)}$. Then, update each particle’s velocity $\mathbf{v}_{n}^{(t)}$ and position $\mathbf{x}_{n}^{(t)}$ by taking into account both $\mathbf{P}_{\text{best},n}^{(t-1)}$ and $\mathbf{G}_{\text{rand}}^{(t-1)}$. If the updated position $\mathbf{x}_{n}^{(t)}$ violates the predefined bounds of the tunable capacitors, a large penalty value is assigned to its fitness $\Gamma_{n}^{(t)}$ to discourage infeasible solutions. After this update, refresh the personal best positions $\mathbf{P}_{\text{best},n}^{(t)}$ and their corresponding reflection coefficients $\Gamma_{\text{best},n}^{(t)}$, as well as the global best position $\mathbf{G}_{\text{best}}^{(t)}$ and its corresponding coefficient $\Gamma_{\text{best}}^{(t)}$. Finally, decrease the annealing temperature $T$.
@@
SAPSO is a non-gradient numerical optimization method with the advantage of high accuracy. It updates parameters by evaluating the fitness of all particles in each iteration. Since computing each particle’s fitness requires invoking the RECBM-Net to infer the corresponding S-parameters, it is computationally intensive. Algorithm 1 provides the detailed procedure for determine the matching solution based on SAPSO.
Classic strategy 2 (AD-Adam): To reduce the number of RECBM-Net evaluations, we adopt adaptive moment optimization estimation (Adam) algorithm [42], a gradient-based method that efficiently updates parameters using first and second moment estimates of gradients. Adam adaptively adjusts the learning rate of each parameter, thereby achieving faster convergence. The effectiveness of Adam critically depends on accurate gradient computation. Automatic differentiation (AD) [43] enables the exact computation of gradients for complex nonlinear functions by constructing computational graphs and systematically applying the chain rule. It overcomes the limitations of both symbolic and numerical differentiation. Therefore, AD is used to compute the gradient of the objective function $\psi\bigl{(}\hat{\mathcal{F}}(f,C_{p},C_{s}),\hat{\Gamma}_{L}\bigr{)}$, which becomes highly nonlinear due to the embedding of the RECBM-Net. The main steps to determine the optimal matching solution using the AD-Adam are as follows.
  • Initialization: Set the learning rate $\alpha$ and exponential decay rates $\beta_{1}$, $\beta_{2}$. Initialize the matching solution as $\mathbf{\theta}_{0}=(C_{p}^{(0)},C_{s}^{(0)})$. Additionally, initialize the first-order moment estimate $\mathbf{m}_{0}$ (momentum term) and the second-order moment estimate $\mathbf{v}_{0}$ for the initial gradient $\mathbf{g}_{0}$.
@@
Proposed strategy (IMS-Net): The two numerical optimization methods discussed above necessitate repeated invocations of RECBM-Net for S-parameters inference, which incurs significant computational overhead. To reduce this burden, we propose a novel strategy that trains an IMS-Net to directly predict the matching solution without iterative DNN evaluation.
Given a predicted $\hat{\Gamma}_{L}$ and $f$ in the current mismatch scenario, the optimal impedance matching solution $(C_{p}^{*},C_{s}^{*})$ satisfies the implicit equation $\psi\bigl{(}\hat{\mathcal{F}}(f,C_{p}^{*},C_{s}^{*}),\hat{\Gamma}_{L}\bigr{)}=0$, assuming a perfect match. This solution can be obtained in two steps. First, the predicted $\hat{\mathbf{S}}$ that satisfies the matching condition is computed via the inverse mapping $\psi^{-1}$. In other words, we solve the equation
@@
Subsequently, the inverse mapping ^$\hat{\mathcal{F}}^{-1}$ is applied to retrieve the circuit configuration corresponding to the obtained scattering property $\hat{\mathbf{S}}$. Accordingly, the matching solution can be expressed as

### v2

@@ II-C Impedance Matching with Parasitic Effects
Considering the inherent limitations of the manufacturing process, parasitic parameters may be introduced at any location along the RF channel connecting the antenna to the transmitter’s RFFE (e.g., radio frequency integrated circuits or power amplifier modules). As shown in Fig. 2, parasitic effects in the RF channel arise from three key aspects.
  • L-network parasitics: These parasitic effects primarily stem from the manufacturing process of the TMN itself [34]. Specifically, whether employing continuously tunable components (e.g., varactors or MEMS varactors) or digitally tunable components (e.g., MEMS switches, p-type-intrinsic-n-type diodes, or complementary metal-oxide-semiconductor switches), each may introduce additional parasitic capacitance, inductance, and resistance. In the ideal case, the L-network transforms the load impedance $Z_{L}$ into the input impedance $Z_{\text{in}}$ via the function $\mathcal{H}(\cdot)$, as described by Eq. (1). The presence of parasitic effects modifies this transformation to a new mapping $\mathcal{H}^{\prime}(\cdot)$, which maps the intermediate impedance $Z_{1}$ to $Z_{2}=\mathcal{H}^{\prime}(Z_{1};f)$, where $f$ denotes the operating frequency.
  • RFFE interface parasitics: These parasitic effects arise from the physical interconnection between the TMN and the RFFE. During the printed circuit board (PCB) routing, via formation, and soldering processes, connection discontinuities or non-ideal contacts can introduce additional parasitic capacitance, inductance, and resistance. The impedance transformation resulting from these parasitic effects is denoted by $\mathcal{H}_{2}(\cdot)$, which maps the intermediate impedance $Z_{2}$ to the final input impedance $Z^{{}^{\prime}}_{\text{in}}=\mathcal{H}_{2}(Z_{2};f)$.
@@
Above all, due to the presence of parasitic effects, the impedance matching condition for ideal L-network given in Eq. (4) is modified to
@@
Neglecting these parasitic effects will lead to impedance matching failure. Taking parasitic effects into account, the overall impedance transformation is expressed as $(\mathcal{H}_{2}\circ\mathcal{H}^{\prime}\circ\mathcal{H}_{1})(\cdot)$, which is strongly nonlinear and inherently uncertain. The strong nonlinearity arises from the fact that parasitic elements are typically superimposed on the RF channel in a mixture of series and shunt configurations. This complex interconnection results in a high-order rational function form for the input impedance. Furthermore, the inherent uncertainty stems from the unpredictable nature of parasitic parameters. Specifically, the types, values, and positions of parasitic parameters are unknown, making the exact mathematical form of $\mathcal{H}_{1}(\cdot)$, $\mathcal{H}_{2}(\cdot)$ and $\mathcal{H}^{\prime}(\cdot)$ indeterminate. Consequently, the presence of parasitic effects not only complicates the relationship between the tunable components and the input impedance but also introduces inherent uncertainty, thereby significantly increasing the difficulty of adaptive impedance matching.
@@ III Behavioral Modeling of Circuits with Parasitic Effects
In this section, we will first employ multiport network theory to demonstrate that S-parameters can effectively characterize the circuit behavior of a TMN in the presence of parasitic effects. Subsequently, we will introduce RECBM-Net, a deep learning model designed to fit the TMN behavior.
@@ III-A Analyzing TMN Using Multiport Network Theory
Parasitic effects cause the practical TMN to deviate from the ideal L-network, resulting in a complex and uncertain topology. Therefore, it is not feasible to characterize TMN behavior from the perspective of impedance transformation. By modeling the practical TMN as a two-port network and analyzing it using multiport network theory [35], we can explore the circuit behavior of the TMN under parasitic effects without requiring detailed knowledge of its internal structure.
Any passive two-port network representing a practical TMN can be equivalently represented as a cascade of series and shunt arms. As illustrated in Fig. 3, the equivalent cascaded network of a practical L-network comprises $M$ series arms and $N$ shunt arms due to parasitic effects. The transmission (ABCD) matrix of a two-port network with a series arm is given by
@@
If the first arm is shunt (rather than series), the first matrix on the right side of Eq. (9) is set to the identity matrix. Similarly, if the last arm is series (rather than shunt), the last matrix is set to the identity matrix. Consequently, the ABCD matrix of any passive two-port network can be expressed in the general form of Eq. (9), regardless of the specific arrangement of series and shunt arms. Finally, based on the conversion relationship between the ABCD matrix and the S-parameters, the S-parameters of the entire two-port network are given by
@@
where $Z_{0}$ represents the characteristic impedance. The S-parameters describe the reflection and transmission characteristics of $n$-port networks in high-frequency. They effectively characterize the circuit behavior of a practical TMN in the presence of parasitic effects. In addition, the S-parameters depend exclusively on the network’s internal topology and are not influenced by the load impedance.
Eqs. (9) and (10) establish the S-parameters of a practical TMN as a function of the operating frequency $f$ and the tunable capacitors $C_{p}$ and $C_{s}$. We denote this mapping as
@@ III-B DNN based modeling
$\mathcal{F}$ is an extremely complex and highly nonlinear function. According to the universal approximation theorem [36], a multilayer feedforward network with a sufficient number of hidden neurons can approximate any continuous function with arbitrary precision. By training a DNN on the measured S-parameters of the practical TMN, we can obtain an accurate estimate $\hat{\mathcal{F}}$ of the mapping $\mathcal{F}$. The preceding analysis of the fundamental mathematical representation of $\mathcal{F}$ is crucial, as it confirms the existence of a mathematical relationship between $(f,C_{p},C_{s})$ and the S-parameters. Leveraging a data-driven approach, the DNN directly establishes an end-to-end nonlinear mapping. This method fundamentally circumvents the challenge of determining detailed topology while enabling accurate characterization of system behavior under parasitic effects.
The structure of the proposed RECBM-Net is illustrated in Fig. 4. It consists of an input module, a TMN operating state encoder, a circuit behavior decoder, an output module and a residual structure. The input vector is defined as $\mathbf{x}=(f,C_{p},C_{s})\in\mathbb{R}^{3}$. The output of the RECBM-Net is an 8-dimensional vector $\mathbf{y}\in\mathbb{R}^{8}$, where each pair of entries corresponds to the real and imaginary components of one of the S-parameters, i.e., $\Re(S_{11})$, $\Im(S_{11})$, $\Re(S_{12})$, $\Im(S_{12})$, $\Re(S_{21})$, $\Im(S_{21})$, $\Re(S_{22})$, and $\Im(S_{22})$. $\Re(\cdot)$ and $\Im(\cdot)$ denote the real and imaginary parts, respectively. The proposed RECBM-Net employs an encoder–decoder structure with nine hidden layers. In the operating state encoder module, the hidden layer dimensionality is initially increased to capture deeper and more abstract features from the input. Specifically, the network progressively expands the feature space with 64, 128, 256, 512, and 1024 neurons. Through the operating state encoder, the 3-dimensional TMN operating state $(f,C_{p},C_{s})$ is transformed into a 1024-dimensional representation, thereby enabling RECBM-Net to capture the intricate nonlinear relationships between the operating state and the corresponding S-parameters. Subsequently, in the circuit behavior decoder module, the network begins to reduce the dimensionality in the later layers with 512, 256, 128, and 64 neurons, effectively decoding the learned features into a compact representation. The final output layer consists of 8 neurons, directly predicting the real and imaginary parts of the S-parameters.
@@ IV Matching Solution Determination Strategy Based on Surrogate Model
In this section, we first formulate the task of determining the matching solution based on a circuit behavior surrogate model as a mathematical optimization problem. Given the difficulty of solving this problem analytically, we will propose three optimization strategies with varying computational overhead.
@@ IV-A Formulating the Matching Optimization Problem
Figure 5: Overview of the proposed data-driven adaptive impedance matching method. The RECBM-Net is first trained using measured S-parameters of the practical L-network. During operation, it predicts the $\hat{\mathbf{S}}_{\text{now}}$ corresponding to the current TMN operating state. According to Eq. (12), the impedance sensor then predicts the $\hat{\Gamma}_{L}$. Finally, the control unit determines the matching solution based on $f$ and $\hat{\Gamma}_{L}$. To achieve this, we propose three strategies that differ in online computational overhead.
@@
where $\psi(\cdot)$ denotes taking the magnitude of the reflection coefficient given in Eq. (13). Additionally, the forward propagation of RECBM-Net for predicting $\mathbf{\hat{S}}$ is denoted as $\hat{\mathcal{F}}(f,C_{p},C_{s})$. Based on this black-box model of TMN behavior, the impedance matching process can be formulated as the following mathematical optimization problem:
@@
where $C_{p,\min}$ and $C_{p,\max}$ denote the lower and upper limits of $C_{p}$, while $C_{s,\min}$ and $C_{s,\max}$ denote those of $C_{s}$. In the current mismatched scenario, the operating frequency $f$ and the derived load reflection coefficient $\hat{\Gamma}_{L}$ are treated as fixed.
Input: Number of particles NN, individual learning factor $\kappa_{1}$, social learning factor $\kappa_{2}$, cooling factor $\lambda$, maximum iterations $\mathcal{I}_{\max}$, load reflection $\hat{\Gamma}_{\text{L}}$, frequency $f$, threshold $\varepsilon$, tuning range $\mathcal{C}=[C_{p,\min},C_{p,\max}]\times[C_{s,\min},C_{s,\max}]$.
@@
19    $\begin{aligned} \mathbf{v}_{n}^{(t)}&=\chi[\mathbf{v}_{n}^{(t-1)}+\kappa_{1}r_{1}(\mathbf{P}_{\text{best},n}^{(t-1)}-\mathbf{x}_{n}^{(t-1)})\\ &\quad+\kappa_{2}r_{2}(\mathbf{G}_{\text{rand}}^{(t-1)}-\mathbf{x}_{n}^{(t-1)})].\end{aligned}$
@@
22       $\Gamma_{n}^{(t)}=\psi(\hat{\mathcal{F}}(f,\mathbf{x}_{n}^{(t)}),\hat{\Gamma}_{L})+P(\mathbf{x}_{n}^{(t)})$.
@@
27      Updated: $\mathbf{P}_{\text{best},n}^{(t)}$, $\Gamma_{\text{best},n}^{(t)}$$\mathbf{G}_{\text{best}}^{(t)}$ and $\Gamma_{\text{best}}$.
@@
2Initialization: $t=0$, $\boldsymbol{\theta}^{(0)}=\boldsymbol{\theta}_{0}$, $\mathbf{m}_{0}=\mathbf{0}$$\mathbf{v}_{0}=\mathbf{0}$.
@@
10   Evaluate reflection coefficient magnitude: $\Gamma_{\text{in}}^{(t)}=\psi(\hat{\mathcal{F}}(f,\boldsymbol{\theta}^{(t)}),\hat{\Gamma}_{L})$.
@@ IV-B Strategies for Solving the Matching Problem
[…]
Classic strategy 1 (SAPSO): SAPSO combines the fast convergence and high accuracy of particle swarm optimization (PSO) with the ability of simulated annealing (SA) to escape local optima by occasionally accepting worse solutions. This hybridization enhances the global search for an optimal TMN configuration. Using SAPSO to obtain the matching solution involves three major steps.
  • Initialization of the population: Let $n$ index the particles, and $t$ index the iterations. At iteration $t=0$, set each particle’s initial position $\mathbf{x}_{n}^{(0)}$ (representing a candidate TMN configuration) and velocity $\mathbf{v}_{n}^{(0)}$ within the given constraints. For each particle $n$, compute the initial fitness $\Gamma_{n}^{(0)}$ and record both its personal best position $\mathbf{P}_{\text{best},n}^{(0)}$ and the global best position $\mathbf{G}_{\text{best}}^{(0)}$ among all particles. Based on $\mathbf{G}_{\text{best}}^{(0)}$, determine the initial temperature $T$ for the simulated annealing component. Additionally, initialize the compression factor $\chi$ and the cooling factor $\lambda$.
  • Velocity and position updates: At iteration $t$, first calculate the normalized Metropolis acceptance probabilities $p_{n}^{(t-1)}$ for each particle $n$ based on its personal best fitness $\Gamma_{\text{best},n}^{(t-1)}$. $\Gamma_{\text{best},n}^{(t-1)}$ denotes the best fitness value achieved by particle $n$ up to the current iteration. Next, perform a roulette wheel selection using $p_{n}^{(t-1)}$ to obtain a candidate global best position $\mathbf{G}_{\text{rand}}^{(t-1)}$. Then, update each particle’s velocity $\mathbf{v}_{n}^{(t)}$ and position $\mathbf{x}_{n}^{(t)}$ by taking into account both $\mathbf{P}_{\text{best},n}^{(t-1)}$ and $\mathbf{G}_{\text{rand}}^{(t-1)}$. If the updated position $\mathbf{x}_{n}^{(t)}$ violates the predefined bounds of the tunable capacitors, we add a penalty term $P(\mathbf{x}_{n}^{(t)})$ to the predicted input reflection coefficient magnitude $\psi(\hat{\mathcal{F}}(f,\mathbf{x}_{n}^{(t)}),\hat{\Gamma}_{L})$, yielding the penalized fitness $\Gamma_{n}^{(t)}$ to discourage infeasible solutions. Accordingly, the penalty term is implemented using the death penalty function, with the penalty constant fixed at 2000. The rationale for this selection is detailed in Appendix A of Supplementary Material. After this update, refresh the personal best positions $\mathbf{P}_{\text{best},n}^{(t)}$ and their corresponding reflection coefficients $\Gamma_{\text{best},n}^{(t)}$, as well as the global best position $\mathbf{G}_{\text{best}}^{(t)}$ and its corresponding coefficient $\Gamma_{\text{best}}$. Finally, decrease the temperature $T$.
@@
SAPSO is a non-gradient numerical optimization method with the advantage of high accuracy. It updates parameters by evaluating the fitness of all particles in each iteration. Since computing each particle’s fitness requires invoking the RECBM-Net to infer the S-parameters, it is computationally intensive. Algorithm 1 provides the detailed procedure for determining the matching solution based on SAPSO.
Classic strategy 2 (AD-Adam): To reduce the number of RECBM-Net evaluations, we adopt Adam [37], a gradient-based optimization method that efficiently updates parameters using first and second moment estimates of gradients. Adam adaptively adjusts the learning rate of each parameter, thereby achieving faster convergence. The effectiveness of Adam critically depends on accurate gradient computation. AD [38] enables the exact computation of gradients for complex nonlinear functions by constructing computational graphs and systematically applying the chain rule. It overcomes the limitations of both symbolic and numerical differentiation. Therefore, AD is used to compute the gradient of the objective function $\psi\bigl(\hat{\mathcal{F}}(f,C_{p},C_{s}),\hat{\Gamma}_{L}\bigr)$ with respect to $C_{p}$ and $C_{s}$, which becomes highly nonlinear due to the embedding of RECBM-Net. When performing backpropagation to compute the gradients, adjoint variables are propagated backward through the computational graph, and the local gradient of each elementary operator is evaluated once. Consequently, the overall computational overhead of a gradient evaluation is approximately equivalent to that of an additional forward pass. The main steps to determine the optimal matching solution using AD-Adam are as follows.
  • Initialization: Set the learning rate $\alpha$ and the exponential decay rates $\beta_{1}$, $\beta_{2}$. Initialize the matching solution as $\boldsymbol{\theta}_{0}=(C_{p}^{(0)},C_{s}^{(0)})$. Additionally, initialize the first-order moment estimate $\mathbf{m}_{0}$ (momentum term) and the second-order moment estimate $\mathbf{v}_{0}$ for the initial gradient $\mathbf{g}_{0}$.
@@
Proposed strategy (IMS-Net): The two numerical optimization methods discussed above necessitate repeated invocations of RECBM-Net for S-parameter inference, which incurs significant computational overhead. To reduce this burden, we propose a novel strategy that trains IMS-Net to directly predict the matching solution without iterative DNN evaluation.
Given a predicted $\hat{\Gamma}_{L}$ and $f$ in the current mismatched scenario, the optimal impedance matching solution $(C_{p}^{*},C_{s}^{*})$ satisfies the implicit equation $\psi\bigl(\hat{\mathcal{F}}(f,C_{p}^{*},C_{s}^{*}),\hat{\Gamma}_{L}\bigr)=0$, assuming a perfect match. This solution can be obtained in two steps. First, the predicted $\hat{\mathbf{S}}$ that satisfies the matching condition is computed via the inverse mapping $\psi^{-1}$. In other words, we solve the equation
@@
Subsequently, the inverse mapping $\hat{\mathcal{F}}^{-1}$ is applied to retrieve the circuit configuration corresponding to the obtained scattering property $\hat{\mathbf{S}}$. Accordingly, the matching solution can be expressed as